# Design and Analysis of Binaural Signal Matching with Arbitrary Microphone Arrays

### v2

@@ 3.2 BFBR as a Special Case of BSM
    The number of beamformers $D$ is set according to the number of assumed sources in BSM design, $Q$.
@@
    The $d$-th beamformer weights $\mathbf{w}_{d}$ in (6) are set according to the $d$-th column of $\mathbf{W}$ in (16) for $d=1,\ldots,Q$. This defines $y(\theta_{d},\phi_{d})$ in (7) according to (18).
@@
    The scaling factors $\{\alpha_{d}\}_{d=1}^{Q}$ in (5) are all set to one.
@@ 4 Generalization of BSM Design to Arbitrary Sound Fields
The BSM method presented above was designed to reproduce binaural signals for a sound field composed of a finite set of sources with known DOAs. In many applications, however, these assumptions may not hold or source information may not be available, thus limiting the use of the method. In addition, it was assumed that the sound sources are uncorrelated, comprising another limitation of applying BSM in real-life acoustic environments which include reverberation, and where the assumption of uncorrelated sources is usually violated. In light of these limitations, this section provides conditions under which BSM may accurately reproduce binaural signals even for complex acoustic environments.
@@ 4.1 BSM Design Valid for Arbitrary Set of Sources
The BSM method designed according to (14) is analyzed for sound fields composed of an arbitrary set of sound sources, i.e, an arbitrary number of sources with arbitrary DOAs. Note that it is still assumed that these sources are spatially white. This is performed by studying the BSM reproduction error in (12). More specifically, since the right term in (12) is a regularization term, the following analysis includes only the left term, that describes the binaural signal error excluding the noise, and it is thus assumed that the signal-to-noise ratio (SNR) is sufficiently high, i.e. $\sigma_{s}^{2}\gg\sigma_{n}^{2}$. Next, assume that the BSM filters $\mathbf{c}_{\text{BSM}}^{l,r}$ produce a sufficiently small error such that it is effectively zero:
$\left\lVert\mathbf{V}^{T}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}-\mathbf{h}^{l,r}\right\rVert_{2}^{2}=0.$ (19)
@@
$\mathbf{V}^{T}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}=\mathbf{h}^{l,r}.$ (20)
@@
$\mathbf{Y}_{N_{V}}\mathbf{V}^{T}_{\mathbf{nm}}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}=\mathbf{Y}_{N_{H}}\mathbf{h}^{l,r}_{\mathbf{nm}},$ (21)
@@
is a $Q\times(N+1)^{2}$ matrix holding the SH functions $Y_{n}^{m}(\theta,\phi)$ of order $n$ and degree $m$ [32], $\mathbf{V}^{T}_{\mathbf{nm}}$ is an $(N_{V}+1)^{2}\times M$ matrix with the $m$-th column holding the SFT coefficients of the transfer function $v(\mathbf{d}_{m};\theta,\phi)$ up to order $N_{V}$ for $m=1,\ldots,M$, and $\mathbf{h}^{l,r}_{\mathbf{nm}}$ is a vector of size $(N_{H}+1)^{2}$ holding the SFT coefficients of the HRTFs up to order $N_{H}$.
@@
$\mathbf{Y}_{N_{V}}\mathbf{V}^{T}_{\mathbf{nm}}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}=\mathbf{Y}_{N_{V}}\begin{bmatrix}\mathbf{h}^{l,r}_{\mathbf{nm}}\\ \mathbf{0}\end{bmatrix},$ (23)
where $\begin{bmatrix}\mathbf{h}^{l,r}_{\mathbf{nm}}\\ \mathbf{0}\end{bmatrix}$ is a zero-padded version of $\mathbf{h}^{l,r}_{\mathbf{nm}}$ to a total length of $\big{(}N_{V}+1\big{)}^{2}$. Then, omitting matrix $\mathbf{Y}_{N_{V}}$ from (23) can be performed assuming that the following holds:
@@
$\mathbf{V}^{T}_{\mathbf{nm}}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}=\begin{bmatrix}\mathbf{h}^{l,r}_{\mathbf{nm}}\\ \mathbf{0}\end{bmatrix}.$ (25)
@@
$\mathbf{Y}^{\prime}_{N_{V}}\mathbf{V}^{T}_{\mathbf{nm}}[\mathbf{c}_{\text{BSM}}^{l,r}]^{*}=\mathbf{Y}^{\prime}_{N_{V}}\begin{bmatrix}\mathbf{h}^{l,r}_{\mathbf{nm}}\\ \mathbf{0}\end{bmatrix}.$ (26)
@@ 4.2 BSM Design Valid for Arbitrary Covariance of Source Signals
Next, the validity of BSM designed according to (14) is analyzed for source signals with an arbitrary covariance matrix $\mathbf{R_{s}}$ in (11). Once again, excluding the noise from the analysis reduces (11) to:
Based on the assumptions and derivations in the previous section, substituting (20) in (27) leads to the binaural signals’ MSE, $\epsilon^{l,r}$, being zero, thus extending the validity of BSM to arbitrary $\mathbf{R_{s}}$. However, in the case where (20) does not hold, such generalization may not be valid. In this case, incorporating $\mathbf{R_{s}}$ (or its estimate) in the BSM design may reduce the error, but this is out of the scope of this paper and is suggested for future work.
@@ 5 Performance Limitations at High Frequencies and a Perceptually-motivated Extension
$\mathbf{V}^{T}[\mathbf{c}^{l,r}]^{*}\approx\mathbf{h}^{l,r},$ (29)
@@
$\epsilon_{\text{abs}}^{l,r}=\sigma_{s}^{2}\left\lVert|\mathbf{V}^{T}[\mathbf{c}^{l,r}]^{*}|-|\mathbf{h}^{l,r}|\right\rVert_{2}^{2}+\sigma_{n}^{2}\left\lVert[\mathbf{c}^{l,r}]^{*}\right\rVert_{2}^{2}.$ (31)
Then, the objective of BSM at these frequencies is to find $\mathbf{c}^{l,r}$ that minimizes (31), formally written as:
@@
The solution to this problem is sometimes referred to as MagLS [40, 41, 29]. Notice that the difference between (31) and (12) is the absolute values on the terms $\mathbf{V}^{T}[\mathbf{c}^{l,r}]^{*}$ and $\mathbf{h}^{l,r}$, which reduces the complexity of the problem in (32), and therefore could lead to reduced errors. This new objective is non-convex, and thus, a global minimum may be unachievable. Nonetheless, various approaches for finding a local minimum to (31) are given in [40] alongside a theoretical analysis of their quality.
The incorporation of MagLS into the BSM method is motivated by the works in [29, 42], which used MagLS for rendering Ambisonics signals. In this work, it is suggested to incorporate the complex least-squares solution to $\epsilon^{l,r}$ in (13) below a predefined cutoff frequency for which (29) holds, with the MagLS solution to $\epsilon_{\text{abs}}^{l,r}$ in (32) employed above the cutoff. The value of the cutoff frequency can be determined empirically by examining the BSM performance of the specific array being used, but a general guideline is to use MagLS at frequencies where the ILD is more important than the ITD for spatial perception, that is, roughly above 1.51.51.5\,1.5kHz [36, 37, 38, 39].
@@ 6 Compensating for Head Rotations with BSM Reproduction
Finally, notice that both rotation types can occur in the same recording and reproduction session. In this case, both can be compensated for separately by modifying the BSM filters in (14) according to (33) and (34). Compensating for head rotations with these equations can potentially degrade the reproduction quality of BSM, as will be shown later. This is because the modified BSM filters reproduce binaural signals that correspond to ears positions that may be relatively far from the microphone positions [44]. This depends on the degree of rotation and on the array configuration.
@@ 7 Simulation Study of BSM with a Semi-circular Array
This section presents an objective performance study based on simulations, and using measures that quantify the quality of binaural reproduction with BSM. The study employs a semi-circular microphone array mounted on a rigid sphere. The configuration is motivated by, and used as a proxy of, arrays on wearable devices, such as microphones mounted on a pair of glasses or a headset, which are currently of great interest. While focusing on a specific array, the analysis presented here can be incorporated into any other array in order to assess the performance of BSM and BSM-MagLS.
@@ 7.2 BSM Accuracy under Static Conditions
As an initial study of BSM reproduction accuracy, the following normalized error measure was defined:
@@ 7.3 Effective SH Order of the ATFs and the HRTFs
For this purpose, the SFT of the left ear HRTF and the ATF corresponding to microphone $m=1$ were calculated with the 2702 Lebdev sampling points. Next, $b_{99}$ was calculated for frequencies in the range of $[75,10000]\,$Hz with 757575\,75Hz resolution, and is presented in Fig. 3. Notice the rapid increase of the maximal SH order of the HRTF with frequency. This may lead to (30) not being satisfied at high frequencies, such that (29) is an under-determined system, which may explain the relatively large errors in Fig. 2. In addition, at frequencies where $b_{99}$ of the HRTF is larger than that of the ATF, some HRTF components may not be reconstructed with sufficient accuracy, which may further explain the decrease in accuracy. Since the frequency range of accurate BSM reproduction is relatively limited, its MagLS extension, which may potentially increase this range, will be studied next.
Figure 3: Effective SH order according to the measure $b_{99}$ in (39) of left ear HRTF and of the ATF corresponding to the 1-st microphone of the semi-circular array with $M=6$ microphones.
@@ 7.4 MagLS Extension of BSM and Head Rotation Compensation
Figure 4 presents the magnitude errors of the left ear with the original BSM filters $\mathbf{c}_{\text{BSM}}^{l}$ and the BSM-MagLS filters $\mathbf{c}_{\text{MLS-BSM}}^{l}$. Since the left and rights ear errors are very similar, only the left ear errors are presented here. Notice that BSM-MagLS produces only slightly smaller errors compared to the original BSM, such that there is no significant improvement for static recording and reproduction conditions. As a follow-up, these errors are studied next with head rotation compensation.
@@
Assume that during the playback of the binaural signals, the listener rotates his/her head by $\Delta\phi$ degrees, as described in Section 6. This is illustrated in Fig. 6, where the reference head position can be seen in Fig. 6(a), and a head rotation of $\Delta\phi$ degrees is illustrated in Fig. 6(b). In the latter case, the HRTF vector $\mathbf{h}^{l,r}$ in the BSM filters $\mathbf{c}_{\text{BSM}}^{l,r}$ and $\mathbf{c}_{\text{MLS-BSM}}^{l,r}$ should be modified to (33) to enable head tracking. The magnitude errors in this case with a head rotation of $\Delta\phi=30^{\circ}$ are presented in Fig 5(a). First, note that the errors of the original BSM have increased significantly for the left ear, by up to approximately 10 dB for frequencies above 1 kHz, compared to the static reproduction conditions in Fig. 4. The right ear errors have also increased, but less significantly. This may be explained by the position of the recording array relative to the ears following the head rotation, as can be seen in Fig 6(b). This rotation distances the left ear from the array, while the right ear remains relatively close to the array, and thus, estimating the left ear signal is more challenging. In addition, note that the accuracy of BSM-MagLS also degrades following the head rotation, but overall it is much more robust to the head rotation.
Similarly, the errors were calculated for a head rotation of $\Delta\phi=60^{\circ}$, which is even more challenging, and are presented in Fig. 5(b). The original BSM produces the binaural signals with even larger errors, which are above -10 dB for frequencies higher than 800 Hz, and above 2 kHz the errors are approximately 0 dB. However, the BSM-MagLS errors remain relatively stable, compared to the results for $\Delta\phi=30^{\circ}$, and the errors remain below -10 dB for frequencies below 5 kHz. Overall, BSM-MagLS is expected to produce the magnitude of the binaural signals much more accurately than the original BSM, when head rotations are compensated for.

### v3

@@ 3.2 BFBR as a Special Case of BSM
    The number of beamformers $D$ should be equal to the number of assumed sources in BSM design, i.e., $D=Q$.
@@
    The $d$-th beamformer weights $\mathbf{w}_{d}$ in (6) should be equal to the $d$-th column of $\mathbf{W}$ in (16) for $d=1,\ldots,Q$. This defines $y(\theta_{d},\phi_{d})$ in (7) according to (18).
@@
    The scaling factors $\{\alpha_{d}\}_{d=1}^{Q}$ in (7) are all set to one, such that (7) and (17) match.
@@ 4 Generalization of BSM Design to Arbitrary Sound Fields
The BSM method presented above was designed to reproduce binaural signals for a sound field composed of a finite set of sources with known DOAs. In many applications, however, these assumptions may not hold or source information may not be available, thus limiting the use of the method. In addition, it was assumed that the sound sources are uncorrelated, comprising another limitation of applying BSM in real-life acoustic environments which include reverberation, and where the assumption of uncorrelated sources is usually violated. In light of these limitations, this section provides conditions under which BSM may accurately reproduce binaural signals even for complex acoustic environments with sufficiently high signal-to-noise (SNR) ratio. No assumptions on specific array geometry will be used, and thus these conditions can be useful for any arbitrary array. First, the validity of BSM will be shown for sound fields composed of an arbitrary set of sound sources, i.e, an arbitrary number of sources with arbitrary DOAs. Then, the assumption of i.i.d. sound sources will be relaxed where the validity of BSM will be shown for arbitrary covariance of source signals.
@@ 4.1 BSM Design Valid for Arbitrary Set of Sources
To analyze the BSM method for sound fields composed of an arbitrary set of sound sources, the BSM reproduction error in (12) is studied. More specifically, since the right term in (12) is a regularization term, the following analysis includes only the left term, that describes the binaural signal error excluding the noise, and it is thus assumed that the SNR is sufficiently high, i.e. $\sigma_{s}^{2}\gg\sigma_{n}^{2}$. Next, assume that the BSM filters $\mathbf{c}_{\text{BSM}}^{l,r}$ produce a sufficiently small error such that it is effectively zero:
$\left\lVert\mathbf{V}^{H}\mathbf{c}_{\text{BSM}}^{l,r}-[\mathbf{h}^{l,r}]^{*}\right\rVert_{2}^{2}=0.$ (19)
@@
$\mathbf{V}^{H}\mathbf{c}_{\text{BSM}}^{l,r}=[\mathbf{h}^{l,r}]^{*}.$ (20)
@@
$\mathbf{Y}_{N_{V}}\mathbf{V}^{H}_{\mathbf{nm}}\mathbf{c}_{\text{BSM}}^{l,r}=\mathbf{Y}_{N_{H}}[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*},$ (21)
@@
is a $Q\times(N+1)^{2}$ matrix holding the SH functions $Y_{n}^{m}(\theta,\phi)$ of order $n$ and degree $m$ [34], $\mathbf{V}^{H}_{\mathbf{nm}}$ is an $(N_{V}+1)^{2}\times M$ matrix with the $m$-th column holding the SFT coefficients of the complex conjugate of the transfer function $v(\mathbf{d}_{m};\theta,\phi)$ up to order $N_{V}$ for $m=1,\ldots,M$, and $\mathbf{h}^{l,r}_{\mathbf{nm}}$ is a vector of size $(N_{H}+1)^{2}$ holding the SFT coefficients of the HRTFs up to order $N_{H}$.
@@
$\mathbf{Y}_{N_{V}}\mathbf{V}^{H}_{\mathbf{nm}}\mathbf{c}_{\text{BSM}}^{l,r}=\mathbf{Y}_{N_{V}}\begin{bmatrix}[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*}\\ \mathbf{0}\end{bmatrix},$ (23)
where $\begin{bmatrix}[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*}\\ \mathbf{0}\end{bmatrix}$ is a zero-padded version of $[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*}$ to a total length of $\big{(}N_{V}+1\big{)}^{2}$. Then, omitting matrix $\mathbf{Y}_{N_{V}}$ from (23) can be performed assuming that the following holds:
@@
$\mathbf{V}^{H}_{\mathbf{nm}}\mathbf{c}_{\text{BSM}}^{l,r}=\begin{bmatrix}[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*}\\ \mathbf{0}\end{bmatrix}.$ (25)
@@
$\mathbf{Y}^{\prime}_{N_{V}}\mathbf{V}^{H}_{\mathbf{nm}}\mathbf{c}_{\text{BSM}}^{l,r}=\mathbf{Y}^{\prime}_{N_{V}}\begin{bmatrix}[\mathbf{h}^{l,r}_{\mathbf{nm}}]^{*}\\ \mathbf{0}\end{bmatrix}.$ (26)
@@ 4.2 BSM Design Valid for Arbitrary Covariance of Source Signals
Next, the validity of BSM filters designed according to (14) is analyzed for source signals with an arbitrary covariance matrix $\mathbf{R_{s}}$ in (11). Once again, excluding the noise from the analysis reduces (11) to:
Based on the assumptions and derivations in the previous section, substituting (20), which yields $[\mathbf{h}^{l,r}]^{T}=[\mathbf{c}_{\text{BSM}}^{l,r}]^{H}\mathbf{V}$, in (27) leads to the binaural signals’ MSE, $\epsilon^{l,r}$, being zero, thus extending the validity of BSM to arbitrary $\mathbf{R_{s}}$. However, in the case where (20) does not hold, such generalization may not be valid. In this case, incorporating $\mathbf{R_{s}}$ (or its estimate) in the BSM design may reduce the error, but this is out of the scope of this paper and is suggested for future work. Nevertheless, Section 5 will present a perceptually-motivated improvement to BSM at high frequencies, where satisfying (20) is more challenging.
@@ 5 Performance Limitations at High Frequencies and a Perceptually-motivated Extension
$\mathbf{V}^{H}\mathbf{c}^{l,r}\approx[\mathbf{h}^{l,r}]^{*},$ (29)
@@
$\epsilon_{\text{abs}}^{l,r}=\sigma_{s}^{2}\left\lVert|\mathbf{V}^{H}\mathbf{c}^{l,r}|-|[\mathbf{h}^{l,r}]^{*}|\right\rVert_{2}^{2}+\sigma_{n}^{2}\left\lVert\mathbf{c}^{l,r}\right\rVert_{2}^{2}.$ (31)
Thus, with (31), the aim is to match only the magnitude values of the HRTFs, ignoring the phase. This coincides with the understanding that for spatial perception, preserving the magnitude values of the binaural signal could be useful to improve reproduction of ILD. Then, the objective of BSM at these frequencies is to find $\mathbf{c}^{l,r}$ that minimizes (31), formally written as:
@@
The solution to this problem is sometimes referred to as MagLS [43, 44, 29]. Notice that the difference between (31) and (12) is the absolute values on the terms $\mathbf{V}^{H}\mathbf{c}^{l,r}$ and $[\mathbf{h}^{l,r}]^{*}$, which reduces the complexity of the problem in (32), and therefore could lead to reduced errors. This new objective is non-convex, and thus, a global minimum may be unachievable. Nonetheless, various approaches for finding a local minimum to (31) are given in [43] alongside a theoretical analysis of their quality.
The incorporation of MagLS into the BSM method is motivated by the works in [29, 45], which used MagLS for rendering Ambisonics signals. Moreover, the solution proposed here is similar to the eMagLS method proposed in [30]. Similar to [30], it is suggested to incorporate the complex least-squares solution to $\epsilon^{l,r}$ in (13) below a predefined cutoff frequency for which (29) holds, with the MagLS solution to $\epsilon_{\text{abs}}^{l,r}$ in (32) employed above the cutoff. The value of the cutoff frequency can be determined empirically by examining the BSM performance of the specific array being used, and typically it should be roughly above 1.5 kHz where the ILD is more important than the ITD for spatial perception [39, 40, 41, 42].
@@ 6 Compensating for Head Rotations with BSM Reproduction
Finally, notice that both rotation types can occur in the same recording and reproduction session. In this case, both can be compensated for separately by modifying the BSM filters in (14) according to (33) and (34). Compensating for head rotations with these equations can potentially degrade the reproduction quality of BSM, as will be shown later. This is because the modified BSM filters reproduce binaural signals that correspond to ear positions that may be relatively far from the microphone positions [47]. This depends on the degree of rotation and on the array configuration.
@@ 7 Simulation Study of BSM with a Semi-circular Array
Having established the theoretical foundation for BSM performance, this section presents a performance analysis using a microphone array mounted on a rigid sphere with a semi-circle geometry as a case study. While the results are specific to this configuration, the presented evaluation methodology is broadly applicable and thus serves as a template for assessing the performance of BSM and BSM-MagLS with other array geometries. This provides a valuable tool for selecting array geometry during the design phase.
@@ 7.2 BSM Accuracy under Static Conditions
In this subsection, the full least-squares solution of BSM is analyzed, i.e., without its MagLS extension. For this purpose, the following normalized error measure was defined:
@@ 7.3 Effective SH Order of the ATFs and the HRTFs
For this purpose, the SFT of the left ear HRTF and the ATF corresponding to microphone $m=1$ were calculated with the 2702 Lebdev sampling points. Next, $b_{99}$ was calculated for frequencies in the range of $[75,10000]\,$Hz with 757575\,75Hz resolution, and is presented in Fig. 4. The figure also shows $\lceil kr\rceil$ calculated using the array radius, for reference. Notice the rapid increase of the maximal SH order of the HRTF with frequency. This may lead to (30) not being satisfied at high frequencies, such that (29) is an under-determined system, which may explain the relatively large errors in Fig. 3. In addition, at frequencies where $b_{99}$ of the HRTF is larger than that of the ATF, some HRTF components may not be reconstructed with sufficient accuracy, which may further explain the decrease in accuracy. Since the frequency range of accurate BSM reproduction is relatively limited, its MagLS extension, which may potentially increase this range, will be studied next.
Figure 4: Effective SH order according to the measure $b_{99}$ in (39) of left ear HRTF and of the ATF corresponding to the 1-st microphone of the semi-circular array with $M=6$ microphones. In addition, $\lceil kr\rceil$ is shown for reference, calculated using the array radius.
@@ 7.4 MagLS Extension of BSM and Head Rotation Compensation
Figure 5 presents the magnitude errors of the left ear with the original BSM filters $\mathbf{c}_{\text{BSM}}^{l}$ and the BSM-MagLS filters $\mathbf{c}_{\text{BSM-MagLS}}^{l}$. Since the left and rights ear errors are very similar, only the left ear errors are presented here. The decrease in error of BSM-MagLS compared to the original BSM averaged across all studied frequencies is $\approx 4.2\,$dB and 3.9 dB for the left and right ears, respectively. Notice that BSM-MagLS produces smaller errors compared to the original BSM, especially at frequencies in the range [400, 1000] Hz and at higher frequencies. This may be attributed to the fact that BSM-MagLS estimates only the magnitude of the HRTFs, while BSM estimates both the magnitude and the phase. However, recall that phase information may be very important for spatial perception, such that the original BSM may produce binaural signals at the lower frequency range which are perceptually better. A more elaborate perceptual study is presented in Section 8. As a follow-up, these errors are studied next with head rotation compensation.
@@
Assume that the listener’s head is rotated by $\Delta\phi$ degrees with respect to the reference head position during the recording phase, as described in Section 6. This is illustrated in Fig. 6, where the reference head position can be seen in Fig. 6(a), and a head rotation of $\Delta\phi$ degrees is illustrated in Fig. 6(b). In the latter case, the HRTF vector $\mathbf{h}^{l,r}$ in the BSM filters $\mathbf{c}_{\text{BSM}}^{l,r}$ and $\mathbf{c}_{\text{BSM-MagLS}}^{l,r}$ should be modified to (33) to enable head tracking. The magnitude errors in this case with a head rotation of $\Delta\phi=30^{\circ}$ are presented in Fig 7(a). First, note that the errors of the original BSM have increased significantly for the left ear, by up to approximately 10 dB for frequencies above 1 kHz, compared to the static reproduction conditions in Fig. 5. The right ear errors have also increased, but less significantly. This may be explained by the position of the recording array relative to the ears following the head rotation, as can be seen in Fig 6(b). This rotation distances the left ear from the array, while the right ear remains relatively close to the array, and thus, estimating the left ear signal is more challenging. In addition, note that the accuracy of BSM-MagLS also degrades following the head rotation, but overall it is much more robust to the head rotation. The increase in error of the original BSM due to the head rotation in this case averaged across all studied frequencies is $\approx 2\,$dB and 0.4 dB for the left and right ears, respectively. For BSM-MagLS, these errors are $\approx 0.1\,$dB and 0.9 dB for the left and right ears, respectively.
Similarly, the errors were calculated for a head rotation of $\Delta\phi=60^{\circ}$, which is even more challenging, and are presented in Fig. 7(b). The original BSM produces the binaural signals with even larger errors, which are above -10 dB for frequencies higher than 800 Hz, and above 2 kHz the errors are approximately 0 dB. However, the BSM-MagLS errors remain relatively stable, compared to the results for $\Delta\phi=30^{\circ}$, and the errors remain below -10 dB for frequencies below 5 kHz. The increase in error of the original BSM due to the head rotation in this case averaged across all studied frequencies is $\approx 2.7\,$dB and 1 dB for the left and right ears, respectively. For BSM-MagLS, these errors are $\approx 0\,$dB and 0.5 dB for the left and right ears, respectively. Overall, BSM-MagLS is expected to produce the magnitude of the binaural signals much more accurately than the original BSM, when head rotations are compensated for.
[…]
(a) Original
[…]
(b) Head rotation